Circular No. 8028

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 X1 (LINEAR) An object reported as asteroidal by LINEAR (discovery observation

below), and placed on the NEO Confirmation Page due to its unusual motion, has been found to be cometary on CCD images taken around Dec. 7.4 UT by D. T. Durig (Sewanee, TN, 0.3-m reflector; 25′′ coma, m_{1} = 15.5), by P. R. Holvorcem and M. Schwartz (Nogales, AZ, 0.81-m telescope; diffuse with coma diameter ∼ 10′′ and a hint of tail at p.a. ∼ 300^{o}), and by G. Hug, R. Valentine, and B. Leifer (Eskridge, KS, 0.30-m reflector; diffuse).

2002 UT  α_{2000}  δ_{2000}  m_{2}
Dec. 5.44440  10^{h}34^{m}06.^{s}76  +19^{o}48′16′′.6  18.0

The available astrometry and the following preliminary parabolic orbital elements appear on MPEC 2002-X46.

T = 2002 Sept. 6.989 TT  ω = 15.^{o}653
Ω = 329.682 }2000.0
q = 0.26947 AU  i = 16.937

SUPERNOVA 2002ic Further to IAUC 8027, M. Hamuy reports that a spectrogram (range

320–920 nm), taken on Nov. 29.11 UT by J. Maza and M. Phillips with the Baade telescope, shows that SN 2002ic (cf. IAUC 8019) is a type-Ia supernova. The spectrum reveals strong Fe III absorption features and a weak Si II 635.5-nm line, similar to the peculiar SN 1999ac near maximum light; superimposed narrow emission lines yield a redshift of 0.066.

SUPERNOVA 2002ji IN NGC 3655 P. Ruiz-Lapuente, University of Barcelona, reports [on behalf of the

‘International Time Programme on Physics and Cosmology’ at the telescopes of the European Northern Observatory, and of the ERTN (cf. IAUC 7987)] that a spectrogram (range 335–905 nm; resolution 1.6 nm) of SN 2002ji (cf. IAUC 8025) was obtained with the Nordic Optical Telescope (+ ALFOSC) at La Palma on Dec. 5.24 UT by S. Benetti, A. Balastegui, S. Basa, D. Guide, J. Mendez, J. Raux, and M. Turatto. The spectrum is that of a type-Ib/c supernova near maximum light, with broad lines of Ca II (H and K), Fe II, and Na I with P-Cyg profiles overimposed on a blue continuum. The expansion velocity (from the Ca II lines) is ∼ 15900 km/s.

2002 December 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green